# references.bib

### v3

@@ 4.2. Proof of the cohomological integrality theorem
and it therefore suffices to prove that the ideal $J_{\lambda}\subseteq\mathcal{A}_{\lambda}^{\mathrm{prim}}$ has finite codimension. We see $\mathcal{A}_{\lambda}^{\mathrm{prim}}$ as the ring of regular functions on the affine algebraic variety $\mathfrak{t}/\mathfrak{g}_{\lambda}$. We prove that the closed subscheme defined by $J_{\lambda}$ is supported on $\{0\}$. If the ideal $J_{\lambda}$ vanishes on $y\in(\mathfrak{t}/\mathfrak{g}_{\lambda})_{\overline{\mathbf{Q}}}\smallsetminus\{0\}$ we let $y^{\prime}\in\mathfrak{t}$ be a lift of $y$ and $\nu$ be given by Lemma 4.17. It satisfies $\mathcal{W}^{\nu=0}(V)=\{\alpha\in\mathcal{W}(V)\mid\langle y,\alpha\rangle=0\}\subseteq\mathcal{W}(V)$ and $\mathcal{W}^{\nu=0}(\mathfrak{g})=\{\alpha\in\mathcal{W}(\mathfrak{g})\mid\langle y,\alpha\rangle=0\}\subseteq\mathcal{W}(\mathfrak{g})$. Moreover, at least one of these inequalities is strict. Indeed, we have $y^{\prime}\not\in\mathfrak{g}_{\lambda}$, which means that either $y^{\prime}$ is not in the center of $\mathrm{Lie}(G^{\lambda})$ and so $\nu$ defines a strict Levi of $G^{\lambda}$ or $y^{\prime}$ is not in the Lie algebra of the kernel of the action $G^{\lambda}\rightarrow\mathrm{GL}(V^{\lambda})$ in which case $V^{\nu}\subsetneq V^{\lambda}$. Therefore, $\nu\prec\lambda=0$, i.e $\overline{\nu}$ is nontrivial. By definition, we have $\prod_{\alpha\in\mathcal{W}^{\nu>0}(V^{\lambda})}\alpha\in J_{\lambda}$.
We also have $\prod_{\alpha\in\mathcal{W}^{\nu>0}(V^{\lambda})}\langle y,\alpha\rangle\neq 0$ by choice of $\nu$, for any $\alpha\in\mathcal{W}^{\nu>0}(V^{\lambda})$, $\langle\nu,\alpha\rangle=0$ if and only if $\langle y,\alpha\rangle=0$. This is a contradiction since we found an element $\prod_{\alpha\in\mathcal{W}^{\nu>0}(V^{\lambda})}\alpha\neq 0$ of $J_{\lambda}$ not vanishing on $y$. ∎
Lemma 4.19.
Let $G_{\lambda}^{0}$ be the neutral component of $G_{\lambda}$. Let $k=\dim\mathfrak{g}_{\lambda}$. Then, $G_{\lambda}^{0}\cong(\mathbf{C}^{*})^{k}$ for some $k\geqslant 0$ and the natural pullback map $\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\rightarrow\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda}^{0})$ is an isomorphism.
Proof.
The group $G_{\lambda}$ is a subgroup of the center of $G^{\lambda}$, and so it is a diagonalisable algebraic group. Therefore, its neutral component is a torus. We have a splitting $G\cong G_{\lambda}^{0}\times\pi_{0}(G_{\lambda})$ [humphreys2012linear, Theorem p.104] and so, by the Künneth formula,
$\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\cong\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda}^{0})\otimes\mathrm{H}^{*}(\mathrm{pt}/\pi_{0}(G_{\lambda})).$
@@
The elements $\alpha\in\mathcal{W}^{\lambda>0}(V)\cup\mathcal{W}^{\lambda>0}(\mathfrak{g})$ (or more generally, $\alpha\in X^{*}(T)$ such that $\langle\lambda,\alpha\rangle\neq 0$), seen as elements of $\mathcal{A}_{\lambda}$, are not zero divisors in the quotient $\mathcal{A}_{\lambda}$-module $\mathcal{A}_{\lambda}^{\mathrm{loc}}/\widetilde{J}_{\lambda}$.
@@
Let $\alpha\in\mathcal{W}^{\lambda>0}(V)\cup\mathcal{W}^{\lambda>0}(\mathfrak{g})$. We see $\alpha\in\mathfrak{t}^{*}$ and $\langle\lambda,\alpha\rangle\neq 0$ implies that $\alpha$ does not vanish on $\mathfrak{g}_{\lambda}$ since it takes a nonzero value on $\left(\frac{\mathrm{d}}{\mathrm{d}t}\lambda(t)\right)_{t=0}\in\mathfrak{g}_{\lambda}$. Therefore, the projection $\alpha^{\prime}\in\mathfrak{g}_{\lambda}^{*}$ of $\alpha$ with respect to the chosen (or any) decomposition $\mathfrak{t}^{*}\cong(\mathfrak{t}/\mathfrak{g}_{\lambda})^{*}\oplus\mathfrak{g}_{\lambda}^{*}$ is nonzero.
We choose a basis $(x_{1},\ldots,x_{\dim\mathfrak{g}_{\lambda}})$ of $\mathfrak{g}_{\lambda}^{*}$ such that $x_{1}=\alpha^{\prime}$ and we order the monomials by the lexicographic order so that $x_{1}\geqslant x_{2}\geqslant\ldots$. Let $g\in\mathcal{A}_{\lambda}^{\mathrm{loc}}\cong\mathcal{A}_{\lambda}^{\mathrm{prim},\mathrm{loc}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$. It may be written in a unique way (see also Lemma 4.19)
@@
Assume that $g\not\in\widetilde{J}_{\lambda}$. We need to show that for $\alpha\in\mathcal{W}^{\lambda>0}(V)\cup\mathcal{W}^{\lambda>0}(\mathfrak{g})$, $\alpha g\not\in\widetilde{J}_{\lambda}$. We let $\nu=\max\{\nu^{\prime}\colon g_{\nu^{\prime}}\neq 0\}$. We may assume that $g_{\nu^{\prime}}\not\in J_{\lambda}$. We have
@@
and $\alpha-\alpha^{\prime}\in\mathrm{Sym}((\mathfrak{t}/\mathfrak{g}_{\lambda})^{*})$. Therefore, the term of highest degree in the variables $(x_{1},\ldots,x_{\dim\mathfrak{g}_{\lambda}})$ in $\alpha g$ is
$g_{\nu}x_{1}x^{\nu}$
and $g_{\nu}\not\in J_{\lambda}$. By the equivalence above, $\alpha g\not\in\widetilde{J}_{\lambda}$. This concludes. ∎
We let $\mathcal{A}_{\lambda}^{\prime}\coloneqq\mathcal{A}_{\lambda}[\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V)\cup\mathcal{W}^{\lambda>0}(\mathfrak{g})}\alpha^{-1}]$, $\mathcal{A}^{\prime\mathrm{loc}}\coloneqq\mathcal{A}_{\lambda}[\alpha^{-1}\colon\alpha\in\mathcal{W}(\mathfrak{g})\smallsetminus\{0\}]$ and $\mathcal{H}_{\lambda}^{\prime}\coloneqq(\mathcal{A}_{\lambda}^{\prime})^{W^{\lambda}}$. We have the localisation maps
$L\colon\mathcal{A}_{\lambda}\rightarrow\mathcal{A}_{\lambda}^{\prime},\quad L\colon\mathcal{H}_{\lambda}\rightarrow\mathcal{H}_{\lambda}^{\prime}.$
We let We let $\widetilde{J}_{\lambda}^{\prime}\coloneqq\mathcal{A}_{\lambda}^{\prime}L(\widetilde{J}_{\lambda})$ be the localised $\mathcal{A}_{\lambda}^{\prime}$-module inside $\mathcal{A}_{\lambda}^{\prime\mathrm{loc}}$.
@@
This follows from Lemma 4.20 as we localise by a set of nonzero divisors of the quotient $\mathcal{A}_{\lambda}^{\mathrm{loc}}/\widetilde{J}_{\lambda}$ in $\mathcal{A}_{\lambda}$. ∎
@@
$k_{\mu}=k_{\nu,\lambda}\frac{\prod_{\alpha\in\mathcal{W}^{\mu>0}(V)\cap\mathcal{W}^{\lambda\neq 0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu>0}(\mathfrak{g})\cap\mathcal{W}^{\lambda\neq 0}(\mathfrak{g})}\alpha}.$
@@
This is immediate from the definitions. ∎
@@
$k^{\prime}_{\mu,\lambda}\coloneqq\frac{\prod_{\alpha\in\mathcal{W}^{\mu>0}(V)\cap\mathcal{W}^{\lambda\neq 0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu>0}(\mathfrak{g})\cap\mathcal{W}^{\lambda\neq 0}(\mathfrak{g})}\alpha}.$
@@
Let $\lambda,\mu\in X_{*}(T)$ and $w$ in $W$ be such that $w\cdot\mu\neq\lambda$ and $\lambda\not\preceq w\cdot\mu$. Then, $w\cdot k_{\mu}\in\widetilde{J}^{\prime}_{\lambda}$.
@@
Of course, the presence of $W$ is superfluous since $wk_{\mu}=k_{w\cdot\mu}$ and it suffices to prove that for any $\lambda,\mu\in X_{*}(T)$ such that $\lambda\not\preceq\mu$, $k_{\mu}\in\widetilde{J}^{\prime}_{\lambda}$.
@@
$k_{\mu}=k_{\nu,\lambda}\frac{\prod_{\alpha\in\mathcal{W}^{\mu>0}(V)\cap\mathcal{W}^{\lambda\neq 0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu>0}(\mathfrak{g})\cap\mathcal{W}^{\lambda\neq 0}(\mathfrak{g})}\alpha}.$
@@
since $\mathcal{A}\cong\mathcal{A}_{\lambda}$ and $W^{\lambda}\subseteq W$ (Lemma 4.2).
@@
$\mathrm{Res}_{\lambda}\left(\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}(f_{\mu})\right)\equiv k_{\lambda}\cdot\sum_{w\in W_{\lambda}/W^{\lambda}}\varepsilon_{V,\lambda}(w)(w\cdot f_{\lambda})\pmod{\widetilde{J}^{\prime}_{\lambda}}.$
@@
$\mathrm{Res}_{\lambda}\left(\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}(f_{\mu})\right)$ $=\mathrm{Res}_{\lambda}\left(\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\frac{1}{\#W^{\mu}}\sum_{w\in W}(w\cdot f_{\mu})(w\cdot k_{\mu})\right)$
@@
$=\frac{1}{\#W^{\lambda}}k_{\lambda}\cdot\sum_{w\in W_{\lambda}}\varepsilon_{V,\lambda}(w)(w\cdot f_{\lambda})\quad\text{by definition of the sign $\varepsilon_{V,\lambda}$}.$
@@
$\bigoplus_{\widetilde{\mu}\in\mathscr{P}_{V}/W}(\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu}))^{\varepsilon_{V,\mu}}\rightarrow\mathrm{H}^{*}(V/G)$
is injective.
@@
Let $f_{\mu}\in(\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu}))^{\varepsilon_{V,\mu}}$, $\widetilde{\mu}\in\mathscr{P}_{V}/W$.
We let $T\coloneqq\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}(f_{\mu})$. We assume that not all $f_{\mu}$’s are zero, and we have to prove that $T\neq 0$. We let $\widetilde{\lambda}\in\mathscr{P}_{V}/W$ be such that $f_{\lambda}\neq 0$ and for any $\widetilde{\mu}\in\mathscr{P}_{V}/W$ such that $f_{\mu}\neq 0$, for any $w\in W$, $\lambda\not\preceq\mu$ (otherwise, we may replace $\lambda$ by $\mu$, etc.).
@@
$\mathrm{Res}_{\lambda}\left(\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}(f_{\mu})\right)\equiv k_{\lambda}\cdot\sum_{w\in W_{\lambda}/W^{\lambda}}\varepsilon(w)(w\cdot f_{\lambda})\equiv k_{\lambda}\#(W_{\lambda}/W^{\lambda})f_{\lambda}\pmod{\widetilde{J}^{\prime}_{\lambda}}.$
and since $f_{\lambda}\not\in\widetilde{J}^{\prime}_{\lambda}$, since $f_{\lambda}\in\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$, and $k_{\lambda}$ is invertible in $\mathcal{A}_{\lambda}^{\prime}$, $\mathrm{Res}_{\lambda}(T)\neq 0$. Therefore, $T\neq 0$, proving injectivity. ∎
@@
$\bigoplus_{\widetilde{\mu}}(\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu}))^{\varepsilon_{V,\mu}}\xrightarrow{\bigoplus_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}}\mathrm{H}^{*}(V/G)$

### v6

@@ 4.2. Proof of the cohomological integrality theorem
because $\mathcal{H}_{\lambda}^{\mathrm{prim}}\cap J_{\lambda}=J_{\lambda}^{W^{\lambda}}$ (second part of Lemma 4.10). Therefore, it suffices to prove that the ideal $J_{\lambda}\subseteq\mathcal{A}_{\lambda}^{\mathrm{prim}}$ has finite codimension. We see $\mathcal{A}_{\lambda}^{\mathrm{prim}}$ as the ring of regular functions on the affine algebraic variety $\mathfrak{t}/\mathfrak{g}_{\lambda}$. We prove that the closed subscheme defined by $J_{\lambda}$ is supported on $\{0\}$. If the ideal $J_{\lambda}$ vanishes on $y\in(\mathfrak{t}/\mathfrak{g}_{\lambda})_{\overline{\mathbf{Q}}}\smallsetminus\{0\}$ (where $\overline{\mathbf{Q}}$ is the field of algebraic numbers) we let $y^{\prime}\in\mathfrak{t}$ be a lift of $y$ and $\nu$ be given by Lemma 4.15. It satisfies $\mathcal{W}^{\nu=0}(V)=\{\alpha\in\mathcal{W}(V)\mid\langle y,\alpha\rangle=0\}\subseteq\mathcal{W}(V)$ and $\mathcal{W}^{\nu=0}(\mathfrak{g})=\{\alpha\in\mathcal{W}(\mathfrak{g})\mid\langle y,\alpha\rangle=0\}\subseteq\mathcal{W}(\mathfrak{g})$. Moreover, at least one of these inequalities is strict. Indeed, we have $y^{\prime}\not\in\mathfrak{g}_{\lambda}$, which means that either $y^{\prime}$ is not in the center of $\mathrm{Lie}(G^{\lambda})$ and so $\nu$ defines a strict Levi $G^{\nu}=G^{y}$ of $G^{\lambda}$ or $y^{\prime}$ is not in the Lie algebra of the kernel of the action $G^{\lambda}\rightarrow\mathrm{GL}(V^{\lambda})$ in which case $V^{\nu}=V^{y}\subsetneq V^{\lambda}$. Therefore, $\nu\prec\lambda=0$, i.e $\overline{\nu}$ is nontrivial. By definition, we have $\prod_{\alpha\in\mathcal{W}^{\nu<0}(V^{\lambda})}\alpha\in J_{\lambda}$.
We also have $\prod_{\alpha\in\mathcal{W}^{\nu<0}(V^{\lambda})}\langle y,\alpha\rangle\neq 0$: indeed, by choice of $\nu$, for any $\alpha\in\mathcal{W}^{\nu<0}(V^{\lambda})$, $\langle\nu,\alpha\rangle=0$ if and only if $\langle y,\alpha\rangle=0$. This is a contradiction since we found an element $\prod_{\alpha\in\mathcal{W}^{\nu<0}(V^{\lambda})}\alpha\neq 0$ of $J_{\lambda}$ not vanishing on $y$. Therefore, $J_{\lambda}$ defines a subscheme of $\mathfrak{t}/\mathfrak{g}_{\lambda}$ supported on 0. ∎
@@
The elements $\alpha\in\mathcal{W}^{\lambda<0}(V)\cup\mathcal{W}^{\lambda<0}(\mathfrak{g})$ (or more generally, $\alpha\in X^{*}(T)$ such that $\langle\lambda,\alpha\rangle\neq 0$), seen as elements of $\mathcal{A}_{\lambda}$, are not zero divisors in the quotient $\mathcal{A}_{\lambda}$-module $\mathcal{A}_{\lambda}^{\mathrm{loc}}/\widetilde{J}_{\lambda}$.
@@
Let $\alpha\in\mathcal{W}^{\lambda<0}(V)\cup\mathcal{W}^{\lambda<0}(\mathfrak{g})$. We see $\alpha\in\mathfrak{t}^{*}$ via the natural inclusion $X^{*}(T)\subseteq\mathfrak{t}^{*}$ and $\langle\lambda,\alpha\rangle\neq 0$ implies that $\alpha$ does not vanish on $\mathfrak{g}_{\lambda}$ since it takes a nonzero value on $\left(\frac{\mathrm{d}}{\mathrm{d}t}\lambda(t)\right)_{t=1}\in\mathfrak{g}_{\lambda}\subseteq\mathfrak{t}$. Therefore, the projection $\alpha^{\prime}\in\mathfrak{g}_{\lambda}^{*}$ of $\alpha$ with respect to the chosen (or any) decomposition $\mathfrak{t}^{*}\cong(\mathfrak{t}/\mathfrak{g}_{\lambda})^{*}\oplus\mathfrak{g}_{\lambda}^{*}$ is nonzero.
We choose a basis $(x_{1},\ldots,x_{\dim\mathfrak{g}_{\lambda}})$ of $\mathfrak{g}_{\lambda}^{*}$ such that $x_{1}=\alpha^{\prime}$ and we order the monomials by the lexicographic order so that $x_{1}\geqslant x_{2}\geqslant\ldots$. Let $g\in\mathcal{A}_{\lambda}^{\mathrm{loc}}\cong\mathcal{A}_{\lambda}^{\mathrm{prim},\mathrm{loc}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$. It may be written in a unique way (see Lemma 4.1)
@@
Assume that $g\not\in\widetilde{J}_{\lambda}$. We need to show that for $\alpha\in\mathcal{W}^{\lambda<0}(V)\cup\mathcal{W}^{\lambda<0}(\mathfrak{g})$, $\alpha g\not\in\widetilde{J}_{\lambda}$. We let $\nu=\max\{\nu^{\prime}\in\mathbf{N}^{\dim\mathfrak{g}_{\lambda}}\colon g_{\nu^{\prime}}\neq 0\}$ (the maximum is taken with respect to the lexicographic order). We may assume that $g_{\nu}\not\in J_{\lambda}$. We have
@@
$=(\alpha-\alpha^{\prime})g+x_{1}g$
and $\alpha-\alpha^{\prime}\in(\mathfrak{t}/\mathfrak{g}_{\lambda})^{*}\subseteq\mathrm{Sym}((\mathfrak{t}/\mathfrak{g}_{\lambda})^{*})$. Therefore, the term of highest degree in the variables $(x_{1},\ldots,x_{\dim\mathfrak{g}_{\lambda}})$ in $\alpha g$ is $g_{\nu}x_{1}x^{\nu}$ and $g_{\nu}\not\in J_{\lambda}$. By the equivalence above, $\alpha g\not\in\widetilde{J}_{\lambda}$. This concludes the proof. ∎
We let $\mathcal{A}_{\lambda}^{\prime}\coloneqq\mathcal{A}_{\lambda}[\prod_{\alpha\in\mathcal{W}^{\lambda<0}(V)\cup\mathcal{W}^{\lambda<0}(\mathfrak{g})}\alpha^{-1}]$, $\mathcal{A}^{\prime\mathrm{loc}}_{\lambda}\coloneqq\mathcal{A}_{\lambda}[\alpha^{-1}\colon\alpha\in\mathcal{W}(\mathfrak{g})\smallsetminus\{0\}]$ and $\mathcal{H}_{\lambda}^{\prime}\coloneqq(\mathcal{A}_{\lambda}^{\prime})^{W^{\lambda}}$. We have the localisation maps
$L\colon\mathcal{A}_{\lambda}\rightarrow\mathcal{A}_{\lambda}^{\prime},\quad L\colon\mathcal{A}_{\lambda}^{\mathrm{loc}}\rightarrow\mathcal{A}_{\lambda}^{\prime\mathrm{loc}},\quad L\colon\mathcal{H}_{\lambda}\rightarrow\mathcal{H}_{\lambda}^{\prime}\,.$
We let $\widetilde{J}_{\lambda}^{\prime}\coloneqq\mathcal{A}_{\lambda}^{\prime}L(\widetilde{J}_{\lambda})$ be the localised $\mathcal{A}_{\lambda}^{\prime}$-module inside $\mathcal{A}_{\lambda}^{\prime\mathrm{loc}}$ (i.e. the $\mathcal{A}_{\lambda}^{\prime}$-submodule of $\mathcal{A}_{\lambda}^{\prime\mathrm{loc}}$ generated by $J_{\lambda}$). We emphasized here the fact that $\widetilde{J}^{\prime}_{\lambda}$ is not a $\mathcal{A}^{\prime\mathrm{loc}}_{\lambda}$-submodule.
@@
This follows from Lemma 4.17 as we localise by a set of nonzero divisors of the quotient $\mathcal{A}_{\lambda}$-module $\mathcal{A}_{\lambda}^{\mathrm{loc}}/\widetilde{J}_{\lambda}$ in $\mathcal{A}_{\lambda}$. More precisely, the first localization morphism $L\colon\mathcal{A}_{\lambda}^{\mathrm{loc}}/\widetilde{J}_{\lambda}\rightarrow\mathcal{A}_{\lambda}^{\prime\mathrm{loc}}/\widetilde{J}_{\lambda}^{\prime}$ is the localization with respect to $\prod_{\alpha\in\mathcal{W}^{\lambda<0}(V)\cup\mathcal{W}^{\lambda<0}(\mathfrak{g})}\alpha$ for the $\mathcal{A}_{\lambda}$-module $\mathcal{A}_{\lambda}^{\mathrm{loc}}/\widetilde{J}_{\lambda}$, and the localization element is not a zero-divisor of $\mathcal{A}_{\lambda}^{\mathrm{loc}}/\widetilde{J}_{\lambda}$. Therefore, the first localization morphism is injective.
Now, the two localization morphisms of the lemma fit into the commutative diagram
(4.2) $\mathcal{H}_{\lambda}/(\widetilde{J}_{\lambda})^{W^{\lambda}}$$\mathcal{H}^{\prime}_{\lambda}/(\widetilde{J}^{\prime}_{\lambda})^{W^{\lambda}}$$\mathcal{A}_{\lambda}^{\mathrm{loc}}/\widetilde{J}_{\lambda}$$\mathcal{A}^{\prime\mathrm{loc}}_{\lambda}/\widetilde{J}^{\prime}_{\lambda}$$\scriptstyle{L}$$\scriptstyle{a}$$\scriptstyle{L}$
The vertical map $a$ is induced by the inclusion $\mathcal{H}_{\lambda}=\mathcal{H}_{\lambda}^{W^{\lambda}}\subseteq\mathcal{A}_{\lambda}\subseteq\mathcal{A}_{\lambda}^{\mathrm{loc}}$. The map $a$ is injective since $\mathcal{H}_{\lambda}\cap\widetilde{J}_{\lambda}=(\widetilde{J}_{\lambda})^{W^{\lambda}}$, which follows from the arguments of the proof of Lemma 4.10. Therefore, the second localization map (the top line of (4.2)) must be injective. ∎
@@
$k_{\mu}=k_{\nu,\lambda}\frac{\prod_{\alpha\in\mathcal{W}^{\mu<0}(V)\cap\mathcal{W}^{\lambda\neq 0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu<0}(\mathfrak{g})\cap\mathcal{W}^{\lambda\neq 0}(\mathfrak{g})}\alpha}\,.$
@@
By the properties satisfied by the cocharacter $\nu$ given by Lemma 2.8, we have $\mathcal{W}^{\mu<0}(V)=\mathcal{W}^{\nu<0}(V)\cap\mathcal{W}^{\lambda=0}(V)\sqcup\mathcal{W}^{\mu<0}(V)\cap\mathcal{W}^{\lambda\neq 0}(V)$ and similarly for $V$ and $\mathfrak{g}$ exchanged. Therefore,
$k_{\mu}$ $=\frac{\prod_{\alpha\in\mathcal{W}^{\mu<0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu<0}(\mathfrak{g})}\alpha}$
$=\frac{\prod_{\alpha\in\mathcal{W}^{\nu<0}(V)\cap\mathcal{W}^{\lambda=0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\nu<0}(\mathfrak{g})\cap\mathcal{W}^{\lambda=0}(\mathfrak{g})}\alpha}\frac{\prod_{\alpha\in\mathcal{W}^{\mu<0}(V)\cap\mathcal{W}^{\lambda\neq 0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu<0}(\mathfrak{g})\cap\mathcal{W}^{\lambda\neq 0}(\mathfrak{g})}\alpha}$
$=k_{\nu,\lambda}\frac{\prod_{\alpha\in\mathcal{W}^{\mu<0}(V)\cap\mathcal{W}^{\lambda\neq 0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu<0}(\mathfrak{g})\cap\mathcal{W}^{\lambda\neq 0}(\mathfrak{g})}\alpha}\,.$
@@
$k^{\prime}_{\mu,\lambda}\coloneqq\frac{\prod_{\alpha\in\mathcal{W}^{\mu<0}(V)\cap\mathcal{W}^{\lambda\neq 0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu<0}(\mathfrak{g})\cap\mathcal{W}^{\lambda\neq 0}(\mathfrak{g})}\alpha}\,.$
@@
Let $\lambda,\mu\in X_{*}(T)$ and $w$ in $W$ be such that $\lambda\not\preceq w^{-1}\cdot\mu$. Then, $w\cdot k_{\mu}\in\widetilde{J}^{\prime}_{\lambda}$.
@@
Of course, the presence of $W$ is superfluous since $wk_{\mu}=k_{w^{-1}\cdot\mu}$ for any $w\in W$ and it suffices to prove that for any $\lambda,\mu\in X_{*}(T)$ such that $\lambda\not\preceq\mu$, $k_{\mu}\in\widetilde{J}^{\prime}_{\lambda}$.
@@
$k_{\mu}=k_{\nu,\lambda}\underbrace{\frac{\prod_{\alpha\in\mathcal{W}^{\mu<0}(V)\cap\mathcal{W}^{\lambda\neq 0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu<0}(\mathfrak{g})\cap\mathcal{W}^{\lambda\neq 0}(\mathfrak{g})}\alpha}}_{=k^{\prime}_{\mu,\lambda}\in\mathcal{A}^{\prime}_{\lambda}}\,.$
@@
since $\mathcal{A}\cong\mathcal{A}_{\lambda}=\mathrm{H}^{*}(\mathrm{pt}/T)$ and $W^{\lambda}\subseteq W$ (Lemma 4.2).
@@
$L\mathrm{Res}_{\lambda}\left(\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}(f_{\mu})\right)\equiv k_{\lambda}\cdot\sum_{w\in W_{\lambda}/W^{\lambda}}\varepsilon_{V,\lambda}(w)^{-1}(w\cdot f_{\lambda})\pmod{\widetilde{J}^{\prime}_{\lambda}}\,$.
@@
[…]
@@
$\bigoplus_{\widetilde{\mu}\in\mathscr{P}_{V}/W}(\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu}))^{\varepsilon_{V,\mu}}\xrightarrow{\bigoplus\mathrm{Ind}_{\mu}}\mathrm{H}^{*}(V/G)$
induced by the parabolic induction morphisms is injective.
@@
Let $f_{\mu}\in(\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu}))^{\varepsilon_{V,\mu}}$, $\widetilde{\mu}\in\mathscr{P}_{V}/W$. We let $U\coloneqq\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}(f_{\mu})$. We assume that not all of the $f_{\mu}$’s are zero, and we have to prove that $U\neq 0$. We let $\widetilde{\lambda}\in\mathscr{P}_{V}/W$ be such that $f_{\lambda}\neq 0$ and for any $\widetilde{\mu}\in\mathscr{P}_{V}/W$ such that $f_{\mu}\neq 0$, for any $w\in W$, $\lambda\not\preceq w\cdot\mu$ (otherwise, we may replace $\widetilde{\lambda}$ by $\widetilde{\mu}$, etc.).
@@
$L\mathrm{Res}_{\lambda}\left(\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}(f_{\mu})\right)$ $\equiv k_{\lambda}\cdot\sum_{w\in W_{\lambda}/W^{\lambda}}\varepsilon_{V,\lambda}(w)^{-1}(w\cdot f_{\lambda})\pmod{\widetilde{J}^{\prime}_{\lambda}}$
[…]
and moreover, $f_{\lambda}\not\in\widetilde{J}^{\prime}_{\lambda}$ since $f_{\lambda}\in\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$, and $k_{\lambda}$ is invertible in $\mathcal{A}_{\lambda}^{\prime}$. We obtain that $\mathrm{Res}_{\lambda}(U)\neq 0$. Therefore, $U\neq 0$, proving injectivity. ∎
We are now ready to prove Theorem 1.1.
@@
$\bigoplus_{\widetilde{\mu}\in\mathscr{P}_{V}/W}(\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu}))^{\varepsilon_{V,\mu}}\xrightarrow{\bigoplus_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}}\mathrm{H}^{*}(V/G)$